what are the steps to solving 8-1/7x=9
Answer by stanbon(75887)

8-(1/7)x=9
----
Add (1/7)x to both sides to get rid of the minus sign:
8 = (1/7)x + 9
---
Subtract 9 from both sides to get:
(1/7)x = 8-9
---
(1/7)x = -1
----
Multiply both sides by 7 to get:
x = -7
==============
Cheers,
Stan H.